OpenStudy (anonymous):

ANSWER PLEASE If 18 meters of high-resistance wire with a diameter of 1.2 mm has a resistance of 10 ohms, what is the resistance of 27 m of the same type of wire with a diameter of 1.5 mm?

OpenStudy (anonymous):

\[R = \rho \frac{ l }{ A }\]

OpenStudy (anonymous):

if 18 m of wire with CSA of 1.16x10^-6 m^2 has a resistance of 10 ohms, calculate rho for that wire...

OpenStudy (anonymous):

\[\huge \frac{ R _{1} }{R _{2} } =\frac{ \rho }{\rho } \frac{ \frac{ l _{1} }{A _{1} } }{ \frac{ l _{2} }{ A _{2} } }\]

OpenStudy (anonymous):

so... how many ohms is that?...

OpenStudy (anonymous):

calc it now

OpenStudy (anonymous):

whelp, equation editor just broke...

OpenStudy (anonymous):

bbl

OpenStudy (anonymous):

cuz i got 1.16 x 10^-6m^2

OpenStudy (anonymous):

for?

OpenStudy (anonymous):

the multiplication thingy earlier

OpenStudy (anonymous):

rho?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

do u have the answer cuz i really need it...